Greg Monfette and Brandon Nunn – The Shape of a Baseball Career and Steroids

I think both your in-class presentation and the final paper turned out very well! The paper is well-written (see the hardcopy for some comments about a few remaining “fine points” that might still be improved a bit). I think the work you did to edit and improve the first draft you showed me really “paid off!” The final product is well-organized and follows the suggested organization from the assignment description very closely. You are using all the terms correctly and with insight. You are clearly on top of your material. This is a first-rate report on your statistical work.

Some Specific Comments

1. The formula you give here for the “variance” V for the Mann-Kendall computation is not completely general. The correction term after the N(N − 1)(2N + 5)/18 would be different depending on how many different groups of repeated values there are, and on how many times each repeated value occurs. For instance, if there were three groups of repeated values all together, one with 3 repeats, one with 4, and one with 2, then the variance would be computed like this:

V = n(n − 1)(2n + 5)/18 − (3)(2)(11)/18 − (4)(3)(13)/18 − (2)(1)(5)/18.

Similarly, your formula for the Mann-Kendall Z is only valid if S > 0, so what you are saying does not cover every possible case. The general computation of the Mann-Kendall Z statistic depends on whether S > 0 or S < 0:

Z = \{ \begin{matrix}\frac{S−1}{\sqrt{V}} & if S > 0 \\ 0 & if S = 0 \\ \frac{S+1}{\sqrt{V}} & if S < 0 \end{matrix}

2. Maybe it would be better to say that Pujols’ data would not be what one would expect to see over a complete career if the player did not use any “chemical enhancements.”

3. I know I suggested that you should not necessarily put in every plot you generated. But since you did discuss Aaron’s career, and also Clemens and Grove in some detail in the text, it would have been nice to have those figures to refer to to back up what you are saying about those players too.

4. A technical point: When you refer to a web site in a references, you should provide a bit more information. Give the full URL for the exact page where you found the information – unless it’s one of the totally unreadable 200-character long ones. Also, provide the date when you accessed the information, since things do tend to change over time (unlike the case for print sources!)

Paper:

Presentation:

Nick Foster and John Vatalaro – Baseball’s Last 50 Years.

On the whole, you have some very interesting ideas here. As I said about the presentation, you have clearly done a lot of work and distilled a number of interesting observations from the data. The idea of comparing the league leaders versus the league averages in these categories is original and enlightening. This is also well-written for the most part (see some comments below on small issues).

However, as a formal written research report, there are a few serious problems with what you have submitted. First, you did not follow the directions about documenting all of your sources. You need at the very least to list the sources of the data you were using in a References section for a paper like this. Get in the habit of doing this right away because it is expected for this kind of formal writing assignment.

Also, it would have been good to follow the suggested outline from the assignment sheet about the final projects even more closely. For instance, you could have put in section headings for the Introduction, Methodology, Results, Discussion, and Conclusion sections to give more structure to what you were presenting and give the reader some signposts about where you were in your argument at all times.

Along those lines, you never really discussed the methodology you were using (the Mann-Kendall trend statistic, the Sen estimated trend line and how they work) at all. It would have been good to say what the methods were and how you were applying them (via the MAKESENS spreadsheet).

Some Specific Comments

1. Say “number” here rather than “amount”

2. Same comment!

3. I dont quite understand your point here. Its true that each team plays (approximately) the same number of games each year. But the league average number of wins per pitcher (say for starting pitchers) would be something meaningful, even though there would be lots of variation (i.e. a big SD) due to injuries, promotions/demotions from/to the minor leagues, etc.

4. You should explain what this is (do not assume the reader already knows).

5. Better: “fans love offense, or maybe “fans love the offensive game.”

6. Say “evolution in

7. No hyphen here: “counterparts”

8. No apostrophe here: “trend lines”

9. You have the wrong Red Sox left fielder here the person who batted .301 in 1968 was Carl Yastrzemski. Ted Williams had retired by 1968.

10. Say almost all of

11. There might be a slightly different alternative or supplementary explanation here too batters did not have the full season to get “up to speed and hit their stride either

Paper:

Presentation: